# Robust Dexterous Grasping of General Objects

### v2

@@ Abstract
The ability to robustly grasp a variety of objects is essential for dexterous robots. In this paper, we present a framework for zero-shot dynamic dexterous grasping using single-view visual inputs, designed to be resilient to various disturbances. Our approach utilizes a hand-centric object shape representation based on dynamic distance vectors between finger joints and object surfaces. This representation captures the local shape around potential contact regions rather than focusing on detailed global object geometry, thereby enhancing generalization to shape variations and uncertainties. To address perception limitations, we integrate a privileged teacher policy with a mixed curriculum learning approach, allowing the student policy to effectively distill grasping capabilities and explore for adaptation to disturbances. Trained in simulation, our method achieves success rates of 97.0% across 247,786 simulated objects and 94.6% across 512 real objects, demonstrating remarkable generalization. Quantitative and qualitative results validate the robustness of our policy against various disturbances.
@@ 3.1 Hand-centric Object Shape Representation
Effectively capturing shape features that generalize for grasping diverse objects remains a fundamental challenge, especially under shape uncertainty arising from limited viewpoints. To address this, we introduce a sparse hand-centric object representation. As shown in Fig. 3, we construct a 51-dimensional vector by concatenating the distance vectors from each finger joint (including the wrist and fingertips) to their nearest points from the object point clouds. This compact representation efficiently encodes local shape features around potential contact regions. Focusing on the interaction-relevant manifold instead of fine-grained global geometry, this representation facilitates generalization to novel object shapes and improves robustness under perceptual uncertainty. Its effectiveness is verified in Sec. 4.4.
@@ B.1 Initial Pose
B.2 Domain Randomization

### v3

@@ Abstract
The ability to robustly grasp a variety of objects is essential for dexterous robots. In this paper, we present a framework for zero-shot dynamic dexterous grasping using single-view visual inputs, designed to be resilient to various disturbances. Our approach utilizes a hand-centric object shape representation based on dynamic distance vectors between finger joints and object surfaces. This representation captures the local shape around potential contact regions rather than focusing on detailed global object geometry, thereby enhancing generalization to shape variations and uncertainties. To address perception limitations, we integrate a privileged teacher policy with a mixed curriculum learning approach, allowing the student policy to effectively distill grasping capabilities and explore for adaptation to disturbances. Trained in simulation, our method achieves success rates of 97.0% across 247,786 simulated objects and 94.6% across 512 real objects, demonstrating remarkable generalization. Quantitative and qualitative results validate the robustness of our policy against various disturbances. Additional results and supplementary videos are provided on our project website.
1ETH Zurich, Switzerland    2 HKUST (Guangzhou), China    3 HKUST, Hong Kong (China)
@@ 3.1 Hand-centric Object Shape Representation
Effectively capturing shape features that generalize for grasping diverse objects remains a fundamental challenge, especially under shape uncertainty arising from limited viewpoints. To address this, we introduce a sparse hand-centric object representation. As shown in Fig. 3, we construct a 51-dimensional vector by concatenating the distance vectors from each finger joint (including the wrist and fingertips) to their nearest points from the object point clouds. This compact representation efficiently encodes interaction-relevant local shape features instead of fine-grained global geometry, which filters out irrelevant shape variations in non-contact areas, facilitates generalization to novel object shapes, and improves robustness under perceptual uncertainty. Its effectiveness is verified in Sec. 4.4.
@@ B.1 Initial Pose
B.2 Object Point Cloud
To extract object point clouds from a continuous visual stream, we first segment the object before grasping using SAM [49], obtaining an initial, unoccluded point cloud. During grasping, we estimate the real-time point cloud under hand occlusion by tracking object pose changes with Cutie [cheng2023putting] and BundleTrack [51]. To mitigate depth noise, particularly for transparent or reflective objects, we apply outlier removal and per-frame smoothing to the initial point cloud over the first 50 frames.
B.3 Action Space
During grasping, our policy predicts both arm and hand target joint angles as actions. When the objects are grasped during deployment, we set fixed target angles for the arm joints to lift the objects.
B.4 Domain Randomization